Given: SSA triangle, A=62°, a=5cm, b=15cm.
Find B, C, c
A.) 48; 70; 22cm
B.) 48; 70; 24.3cm
C.) 48; 70; 26.7cm
D.) Triangle can't be made.

Answer by Theo(13342)

side a is opposite angle A
side b is opposite angle B
side c is opposite angle C


by the law of sines, you get:

a/sin(A) = b/sin(B) = c/sin(C)

you are given a, b, and A.

use these to get:

a/sin(A) = b/sin(B)

since a = 5 and b = 15 and A = 62 degrees, you get:

5/sin(62) = 15/sin(B)

solve for sin(B) to get sin(B) = 15/5 * sin(62)

solve for sin(B) to get sin(B) = 2.648

since sin(B) can't be greater than 1, your triangle can't be made, to the best of my knowledge.